88456:

Jaden and his family are using a map to drive down to Napa Valley. On the map, the distance is given by the equation c = 1.3m, where c is the number of centimeters on the map and m is the corresponding real-world distance in miles. If they cover a distance of 18 miles, that would mean a distance of centimeters on the map.

88456:

there is a blank between of and cm

jhonyy9:

so more clearly - understandably 1 cm on map mean 1,3 miles in reality what is the distance in cm on mape if they cover a distance of 18 miles ? so to calcule this you can use the rule of 3 simple and write in this way 1 cm on mape mean 1,3 miles x cm mean 18 miles - cross multiplie and calcule the x value x = 18*1/1,3 = 18/1,3 = ? cm hope helped understandably easy